Net Income 6958
Payout ratio 40%
Repurchase 300
What is Cash flow to Stockholders?
Answers:

3687.29

2784.25

3083.2

2483.2
Answer #1

CASH FLOW TO STOCKHOLDERS: 3083.2

  • Cash flow to stockholders = Dividend payments + Stock repurchase = 2783.2 + 300 = 3083.2
    • Dividend payments = net income × payout ratio = 6958 × 40% = 2783.2
    • Stock repurchase: company buys back it's shares from the shareholders by paying cash.